A 2 ft vertical post casts a 16 inch shadow at the same time a nearby cell phone tower casts a 120 ft shadow. How tall is the cell phone tower?
Answer by stanbon(75887)

A 2 ft vertical post casts a 16 inch shadow at the same time a nearby cell phone tower casts a 120 ft shadow. How tall is the cell phone tower?
-------------
Use a proportion::
tower/shadow = post/shadow
-----
16 in = 16/12 ft = 4/3 ft
t/120 ft = 2 ft/(4/3)ft
-----
t = 120[6/4] = 180 ft (tower's height)
-----------------
Cheers,
Stan H.
-------------